Chapter 16, Problem 056 Suppose that is the intensity at a spot that is 82 m away? a public address system emits sound uniformly in all directions and that there are no rene tions. The intersty at .locaton 26 m away horn the sand เฉ¡ce as 3.0 x 204 w/m2 . what are no reflections the tolerance is +/-2%
Answer #1

Since the source is same, power by both will be same,

i1*4pi*r1^2 = i2*4pi*r2^2

3e-4*25^2 = i2*82^2

i2 = 3e-4*25^2/82^2 = 0.0000279 W/m^2